OpenStudy (anonymous):

If 6x ÷ 8 = 40, then _____ . (1 point) 6x ÷ 8 = 320 40 = 6x · 8 6x = 320 40 = 6x ÷ 8

OpenStudy (anonymous):

Which one means the same as 6x ÷ 8 = 40. A) 6x ÷ 8 = 320 B) 6x · 8 = 40 C) 6x = 320 D) 6x ÷ 8 = 40

OpenStudy (anonymous):

6x/8=320

OpenStudy (anonymous):

@Blazettequestioner 6x /8 = 40 is not the same as 6x /8 = 320

OpenStudy (anonymous):

Thanks;)

OpenStudy (anonymous):

Welcome (: